# GAICo: A Deployed and Extensible Framework for Evaluating Diverse and Multimodal Generative AI Outputs

### v2

@@ Abstract
The rapid proliferation of Generative AI (GenAI) into diverse, high-stakes domains necessitates robust and reproducible evaluation methods. However, practitioners often resort to ad-hoc, non-standardized scripts, as common metrics are often unsuitable for specialized, structured outputs (e.g., automated plans, time-series) or holistic comparison across modalities (e.g., text, audio, and image). This fragmentation hinders comparability and slows AI system development. To address this challenge, we present GAICo (Generative AI Comparator): a deployed, open-source Python library that streamlines and standardizes GenAI output comparison. GAICo provides a unified, extensible framework supporting a comprehensive suite of reference-based metrics for unstructured text, specialized structured data formats, and multimedia (images, audio). Its architecture features a high-level API for rapid, end-to-end analysis, from multi-model comparison to visualization and reporting, alongside direct metric access for granular control. We demonstrate GAICo’s utility through a detailed case study evaluating and debugging complex, multi-modal AI Travel Assistant pipelines. GAICo empowers AI researchers and developers to efficiently assess system performance, make evaluation reproducible, improve development velocity, and ultimately build more trustworthy AI systems, aligning with the goal of moving faster and safer in AI deployment. Since its release on PyPI in Jun 2025, the tool has been downloaded over 13K times, across versions, by Aug 2025, demonstrating growing community interest.
@@
Documentationai4society.github.io/projects/GenAIResultsComparator
@@ Introduction
The adoption of Generative AI (GenAI), particularly Large Language Models (LLMs), has been transformative across industries. However, this rapid integration has outpaced the development of standardized, accessible, and domain-appropriate evaluation tools. While standard Natural Language Processing (NLP) metrics like BLEU (Papineni et al. 2002) and ROUGE (Lin 2004) are useful, they are often insufficient for structured outputs like AI plans or time-series forecasts. In the absence of a unified tool, developers create bespoke evaluation scripts, leading to a fragmented and non-reproducible landscape that hinders the comparison and iterative improvement of AI systems.
@@
Figure 1: The multi-modal GAICo workflow. The framework processes answers from multi-modal (text, image, audio) AI models, computes pairwise similarity scores ($s_{kl}$), and constructs several outputs: raw data reports, visualizations, and pass/fail assessments against a threshold $\delta$ (any distance function, or conversely, 1 - similarity metric). See metrics in Table 5.
@@
In response, we introduce GAICo (Generative AI Comparator), a deployed and extensible open-source Python library built to address these challenges. GAICo provides a unifying framework that makes the application of diverse, reference-based metrics simple and reproducible across different data types (modalities). By providing a consistent interface, GAICo directly supports the iterative workflows common in applied AI development. According to a PyPI package statistics source (PyPI Stats 2025), the library has been downloaded over 13,000 times in just over 2 months (Jun-Aug 2025), demonstrating its widespread and growing community interest.
@@
General-purpose AI and Natural Language Processing (NLP) evaluation libraries provide foundational metrics for various tasks. For instance, the Hugging Face evaluate library (Wolf et al. 2019) offers a wide array of metrics for NLP and related tasks, providing a standardized compute method for individual metrics and integrating well within the Hugging Face ecosystem. Similarly, Scikit-learn (Pedregosa et al. 2011) provides fundamental metrics for classification, regression, and clustering, adaptable for certain aspects of text evaluation (e.g., cosine similarity on TF-IDF vectors). Libraries like NLTK (Bird, Klein, and Loper 2009) and SpaCy (Honnibal et al. 2020) offer foundational text processing capabilities and some basic metrics (e.g., BLEU in NLTK).
@@
A key differentiator for GAICo is its design choice to decouple evaluation from LLM inference, positioning it as a post-hoc, reproducible comparison framework. This contrasts with general-purpose libraries like Hugging Face evaluate by offering a more integrated workflow and specialized metrics, and with end-to-end frameworks, e.g., Ragas (Es et al. 2024), DeepEval (Ip and Vongthongsri 2025), that are tightly coupled with LLM APIs for “LLM-as-a-judge” evaluations. While these integrated frameworks are powerful, their coupling can introduce complexities like API costs, rate-limits, and nondeterminism. GAICo’s focus on pre-generated outputs simplifies the evaluation pipeline, making it faster and more reliable for developers focused on comparative analysis. The Experiment class, discussed in Section 3The GAICo Framework, further streamlines this by offering a high-level abstraction for multi-model comparison, plotting, and reporting. A detailed comparison of these approaches is provided in Section 1B of the Appendix.
@@
Automated Planning: Evaluation in automated planning typically focuses on properties like plan length, cost, or state-space coverage, often requiring specialized planners or validators (e.g., PDDL parsers). However, when LLMs generate planning sequences, the output might not strictly adhere to formal planning languages, necessitating metrics that compare the content and order of actions rather than strict plan validity. GAICo addresses this with PlanningLCS and PlanningJaccard metrics, which are designed to robustly compare action sequences, including concurrent actions, directly from LLM outputs. These are inspired by foundational work on measuring plan diversity and similarity (Srivastava et al. 2006, 2007) and sequence analysis (Hirschberg 1977), adapted specifically for GenAI evaluation contexts.
Time-Series Analysis: Standard time-series evaluation metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), or R-squared for forecasting accuracy; however, when comparing generated time-series shapes or sequences of values from LLMs, metrics like Dynamic Time Warping (DTW) (Berndt and Clifford 1994) become crucial for handling variations in speed or phase. GAICo integrates TimeSeriesDTW and TimeSeriesElementDiff to provide robust, normalized comparisons of time-series data, regardless of their textual representation.
Multimedia (Image/Audio): Evaluation of generated images and audio often involves specialized libraries and metrics. For images, common metrics include Structural Similarity Index (SSIM) (Wang et al. 2004), Peak Signal-to-Noise Ratio (PSNR) (Sara, Akter, and Uddin 2019), or perceptual hashes (Farid 2021). For audio, metrics like Signal-to-Noise Ratio (SNR) (Yuan et al. 2019), Perceptual Evaluation of Speech Quality (PESQ) (Rix et al. 2001), or Short-Time Objective Intelligibility (STOI) (Taal et al. 2010) are used. While underlying libraries (e.g., scikit-image (Van der Walt et al. 2014), Pillow (Clark 2015), librosa (McFee et al. 2015), scipy (Virtanen et al. 2020)) provide these functionalities, GAICo unifies their application within its framework. The rise of multi-modal generative models, such as Meta’s ImageBind (Girdhar et al. 2023), further underscores the need for unified evaluation tools capable of assessing outputs across diverse modalities. This allows for consistent, normalized comparison of multimedia outputs alongside text and structured data, bridging the gap between general LLM evaluation and highly specialized domain-specific analysis.
@@
GAICo provides a carefully designed library of metrics spanning textual analysis, structured data, and multimedia (audio and image). This comprehensive approach addresses the reality that modern GenAI applications produce complex, multi-modal outputs that traditional text-only tools cannot adequately assess. The library includes n-gram-based (BLEU, ROUGE), text similarity (Jaccard, Cosine), and semantic (BERTScore) metrics. For structured data, it offers specialized metrics for automated planning sequences (PlanningLCS) and time-series (TimeSeriesDTW). For multimedia, it integrates established metrics for image (SSIM, PSNR) and audio (SNR, Spectrogram Distance) quality. Detailed descriptions of each metric are available in Section 3D of the Appendix.
@@
GAICo is a professionally engineered tool deployed on the Python Package Index (PyPI) for easy installation via pip. To balance functionality with accessibility, it uses optional dependencies (e.g., pip install ‘gaico[bertscore]’), minimizing the installation footprint. The project’s quality is ensured through a comprehensive test suite (pytest (Krekel et al. 2004)), continuous integration (CI) for automated testing and deployment, and pre-commit hooks (Sottile 2014) for code consistency. Extensive documentation (MkDocs (Christie 2014)), including runnable Jupyter notebooks and a full API reference, lowers the barrier to adoption for the applied AI community.
@@
Table 1: Plan Coherence Results. Scores are averaged over three days, comparing each pipeline’s orchestrator output against a single baseline reference. Higher scores between Pipeline B and C are highlighted.
Pipeline ROUGE-L BERTScore-F1 PlanningLCS PlanningJaccard TimeSeriesDTW
Pipeline A 1.000 1.000 1.000 1.000 1.000
Pipeline B 0.190 0.599 0.095 0.083 0.122
Pipeline C 0.222 0.613 0.137 0.117 0.367
Table 2: Modality Generation Quality Results. Scores are averaged over three days, comparing each pipeline’s specialist model outputs to a high-quality reference generated from the same prompt. Higher scores between Pipeline B and C are highlighted.
Pipeline Image SSIM Image AvgHash Image HistMatch Audio SNR Audio SpecDist
Pipeline A 1.000 1.000 1.000 1.000 1.000
Pipeline B 0.276 0.646 0.642 0.249 0.261
Pipeline C 0.347 0.766 0.528 0.247 0.260
@@
The quantitative results, summarized in Tables 1 and 2 and visualized in Figure 3, demonstrate GAICo’s value in accelerating AI development and improving system safety.
@@
As shown in Table 1, Pipeline A’s orchestrator (GPT-5), serving as the high-quality baseline, expectedly achieved perfect scores (1.000) across all plan coherence metrics. In contrast, both Pipeline B (Llama 4) and Pipeline C (Gemini 2.5 Pro) showed significantly lower performance, indicating substantial deviations from the baseline plan. As shown in the radar plot in Figure 3 (left), Pipeline C generally outperformed Pipeline B across these metrics, suggesting a better ability to approximate the reference budget forecast.
@@
The modality generation quality results (Table 2) reveal the execution fidelity of the specialist models against their respective high-quality references. Pipeline A’s specialist models achieved perfect scores (1.000) as their outputs were used to define the modality-specific references. For image generation, Pipeline C generally outperformed Pipeline B across Image SSIM and Image AverageHash, indicating stronger structural and perceptual similarity to the baseline. However, Pipeline B showed a slightly higher Image HistogramMatch score, suggesting its color distribution was closer to the baseline. This highlights the value of multi-metric evaluation for a nuanced understanding.
@@
Improved Reliability and Safer Deployment The two-part evaluation, enabled by GAICo, is critical for enhancing system reliability. By separating “Plan Coherence” from “Modality Generation Quality”, we can precisely diagnose failures. For instance, the results show that Pipeline B’s poor performance stems from both a weak orchestrator (low scores in Table 1) and less-than-optimal specialist models (Table 2). This granular insight allows for targeted optimization, improving the orchestrator’s planning logic or replacing a specific specialist model, which is essential for building trustworthy systems and preventing incidents caused by underperforming components. This directly contributes to the goal of producing better AI solutions by enabling a more rigorous and efficient development process.
@@
Table 3: Summary of GAICo Example Notebooks. Those referenced in paper are emphasized. All available at:
@@ Conclusion
GAICo addresses a critical gap in GenAI evaluation by providing the first unified, extensible, and deployed framework for reproducible assessment across diverse data modalities. Recent studies have found that even with access to the original source code, many ML experiments cannot be reproduced, mainly due to missing details in documentation and inconsistent evaluation procedures (Lopresti and Nagy 2021). GAICo replaces this current landscape of ad-hoc, non-comparable scripts with a standardized approach, integrating metrics for text, structured data (planning, time-series), and multimedia. This standardization is GAICo’s core novelty; it ensures that once a metric is chosen, its application and reporting are consistent, directly improving the quality and productivity of the evaluation process. As demonstrated in our case study, this unified approach enables nuanced debugging, such as distinguishing orchestrator faults from specialist model failures, which is crucial for enhancing the reliability of complex AI systems.
@@
We now discuss how one may extend the limitations in the future. A key priority is to incorporate metrics for fairness, bias, toxicity, and operational statistics, providing developers with a more holistic toolkit for building safer and more responsible AI systems. GAICo’s extensible design makes this easy (see UML class diagram, Figure 7). Future work also includes exploring developing interactive dashboards or deeper integration with MLOps platforms (like MLflow (Zaharia et al. 2018)) to enable dynamic exploration and tracking of evaluation results. Lastly, further work would also focus on adding general-purpose comparison capabilities for complex structured data (e.g., JSON diffing, graph similarity) and extending metrics to handle multi-turn conversational outputs.

### v3

@@ Abstract
The rapid proliferation of Generative AI (GenAI) into diverse, high-stakes domains necessitates robust and reproducible evaluation methods. However, practitioners often resort to ad-hoc, non-standardized scripts, as common metrics are often unsuitable for specialized, structured outputs (e.g., automated plans, time-series) or holistic comparison across modalities (e.g., text, audio, and image). This fragmentation hinders comparability and slows AI system development. To address this challenge, we present GAICo (Generative AI Comparator): a deployed, open-source Python library that streamlines and standardizes GenAI output comparison. GAICo provides a unified, extensible framework supporting a comprehensive suite of reference-based metrics for unstructured text, specialized structured data formats, and multimedia (images, audio). Its architecture features a high-level API for rapid, end-to-end analysis, from multi-model comparison to visualization and reporting, alongside direct metric access for granular control. We demonstrate GAICo’s utility through a detailed case study evaluating and debugging complex, multi-modal AI Travel Assistant pipelines. GAICo empowers AI researchers and developers to efficiently assess system performance, make evaluation reproducible, improve development velocity, and ultimately build more trustworthy AI systems, aligning with the goal of moving faster and safer in AI deployment. Since its release on PyPI in Jun 2025, the tool has been downloaded over 15K times, across versions, by Nov 2025, demonstrating growing community interest.
@@
Documentationai4society.github.io/projects/GenAIResultsComparatsor
@@ Introduction
The adoption of Generative AI (GenAI), particularly Large Language Models (LLMs), has been transformative across industries. However, this rapid integration has outpaced the development of standardized, accessible, and domain-appropriate evaluation tools. While standard Natural Language Processing (NLP) metrics like BLEU (papineni-etal-2002-bleu) and ROUGE (lin-2004-rouge) are useful, they are often insufficient for structured outputs like AI plans or time-series forecasts. In the absence of a unified tool, developers create bespoke evaluation scripts, leading to a fragmented and non-reproducible landscape that hinders the comparison and iterative improvement of AI systems.
Figure 1: The multi-modal GAICo workflow. The framework processes answers from multi-modal (text, image, audio) AI models, computes pairwise similarity scores ($s_{kl}$), and constructs several outputs: raw data reports, visualizations, and pass/fail assessments against a threshold $\delta$. (any distance function, or conversely, 1similarity metric1-$\text${similarity metric}).
In response, we introduce GAICo (Generative AI Comparator), a deployed and extensible open-source Python library built to address these challenges. GAICo provides a unifying framework that makes the application of diverse, reference-based metrics simple and reproducible across different data types (modalities). By providing a consistent interface, GAICo directly supports the iterative workflows common in applied AI development. According to a PyPI package statistics source (pepy_gaico), the library has been downloaded over 13,000 times in just over 2 months (Jun-Aug 2025), demonstrating its widespread and growing community interest.
@@
General-purpose AI and Natural Language Processing (NLP) evaluation libraries provide foundational metrics for various tasks. For instance, the Hugging Face evaluate library (wolf2019huggingface) offers a wide array of metrics for NLP and related tasks, providing a standardized compute method for individual metrics and integrating well within the Hugging Face ecosystem. Similarly, Scikit-learn (scikit-learn) provides fundamental metrics for classification, regression, and clustering, adaptable for certain aspects of text evaluation (e.g., cosine similarity on TF-IDF vectors). Libraries like NLTK (bird2009natural) and SpaCy (Honnibal_spaCy_Industrial-strength_Natural_2020) offer foundational text processing capabilities and some basic metrics (e.g., BLEU in NLTK).
@@
A key differentiator for GAICo is its design choice to decouple evaluation from LLM inference, positioning it as a post-hoc, reproducible comparison framework. This contrasts with general-purpose libraries like Hugging Face evaluate by offering a more integrated workflow and specialized metrics, and with end-to-end frameworks, e.g., Ragas (es2024ragas), DeepEval (Ip_deepeval_2025), that are tightly coupled with LLM APIs for “LLM-as-a-judge” evaluations. While these integrated frameworks are powerful, their coupling can introduce complexities like API costs, rate-limits, and nondeterminism. GAICo’s focus on pre-generated outputs simplifies the evaluation pipeline, making it faster and more reliable for developers focused on comparative analysis. The Experiment class, discussed in Section 3The GAICo Framework, further streamlines this by offering a high-level abstraction for multi-model comparison, plotting, and reporting. A detailed comparison of these approaches is provided in Section 1B of the Appendix.
@@
Automated Planning: Evaluation in automated planning typically focuses on properties like plan length, cost, or state-space coverage, often requiring specialized planners or validators (e.g., PDDL parsers). However, when LLMs generate planning sequences, the output might not strictly adhere to formal planning languages, necessitating metrics that compare the content and order of actions rather than strict plan validity. GAICo addresses this with PlanningLCS and PlanningJaccard metrics, which are designed to robustly compare action sequences, including concurrent actions, directly from LLM outputs. These are inspired by foundational work on measuring plan diversity and similarity (srivastava2006finding; srivastava2007domain) and sequence analysis (hirschberg1977algorithms), adapted specifically for GenAI evaluation contexts.
Time-Series Analysis: Standard time-series evaluation metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), or R-squared for forecasting accuracy; however, when comparing generated time-series shapes or sequences of values from LLMs, metrics like Dynamic Time Warping (DTW) (berndt1994using) become crucial for handling variations in speed or phase. GAICo integrates TimeSeriesDTW and TimeSeriesElementDiff to provide robust, normalized comparisons of time-series data, regardless of their textual representation.
Multimedia (Image/Audio): Evaluation of generated images and audio often involves specialized libraries and metrics. For images, common metrics include Structural Similarity Index (SSIM) (wang2004image), Peak Signal-to-Noise Ratio (PSNR) (sara2019image), or perceptual hashes (farid2021overview). For audio, metrics like Signal-to-Noise Ratio (SNR) (yuan2019signal), Perceptual Evaluation of Speech Quality (PESQ) (rix2001perceptual), or Short-Time Objective Intelligibility (STOI) (taal2010short) are used. While underlying libraries (e.g., scikit-image (van2014scikit), Pillow (clark2015pillow), librosa (mcfee2015librosa), scipy (2020SciPy-NMeth)) provide these functionalities, GAICo unifies their application within its framework. The rise of multi-modal generative models, such as Meta’s ImageBind (girdhar2023imagebind), further underscores the need for unified evaluation tools capable of assessing outputs across diverse modalities. This allows for consistent, normalized comparison of multimedia outputs alongside text and structured data, bridging the gap between general LLM evaluation and highly specialized domain-specific analysis.
@@
GAICo provides a carefully designed library of metrics spanning textual analysis, structured data, and multimedia (audio and image). This comprehensive approach addresses the reality that modern GenAI applications produce complex, multi-modal outputs that traditional text-only tools cannot adequately assess. The library includes n-gram-based (BLEU, ROUGE), text similarity (Jaccard, Cosine), and semantic (BERTScore) metrics. For structured data, it offers specialized metrics for automated planning sequences (PlanningLCS) and time-series (TimeSeriesDTW). For multimedia, it integrates established metrics for image (SSIM, PSNR) and audio (SNR, Spectrogram Distance) quality.
@@
GAICo is a professionally engineered tool deployed on the Python Package Index (PyPI) for easy installation via pip. To balance functionality with accessibility, it uses optional dependencies (e.g., pip install ‘gaico[bertscore]’), minimizing the installation footprint. The project’s quality is ensured through a comprehensive test suite (pytest (pytest)), continuous integration (CI) for automated testing and deployment, and pre-commit hooks (precommit_2025) for code consistency. Extensive documentation (MkDocs (mkdocs_2025)), including runnable Jupyter notebooks and a full API reference, lowers the barrier to adoption for the applied AI community.
@@
Plan Coherence Metrics Modality Generation Metrics
Pipeline $\text{ROUGE-L}_{\textit{T}}$ $\text{BERTScore-F1}_{\textit{T}}$ $\text{LCS}_{\textit{Pl}}$ $\text{Jaccard}_{\textit{Pl}}$ $\text{DTW}_{\textit{TS}}$ $\text{SSIM}_{\textit{I}}$ $\text{AvgHash}_{\textit{I}}$ $\text{HistMatch}_{\textit{I}}$ $\text{SNR}_{\textit{A}}$ $\text{SpecDist}_{\textit{A}}$
Pipeline A 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Pipeline B 0.190 0.599 0.095 0.083 0.122 0.276 0.646 0.642 0.249 0.261
Pipeline C 0.222 0.613 0.137 0.117 0.367 0.347 0.766 0.528 0.247 0.260
Table 1: Combined evaluation results. Plan coherence metrics (left) and modality generation metrics (right) are reported for each pipeline. Subscripts denote evaluation domains: T = text, Pl = planning, TS = time series, I = image, and A = audio. Scores are averaged over three days; higher scores between Pipeline B and Pipeline C are highlighted.
@@
The quantitative results, summarized in Table 1 and visualized in Figure 3, demonstrate GAICo’s value in accelerating AI development and improving system safety.
@@
As shown in the left side of Table 1, Pipeline A’s orchestrator (GPT-5), serving as the high-quality baseline, expectedly achieved perfect scores (1.000) across all plan coherence metrics. In contrast, both Pipeline B (Llama 4) and Pipeline C (Gemini 2.5 Pro) showed significantly lower performance, indicating substantial deviations from the baseline plan. As shown in the radar plot in Figure 3 (left), Pipeline C generally outperformed Pipeline B across these metrics, suggesting a better ability to approximate the reference budget forecast.
@@
The modality generation quality results (right of Table 1) reveal the execution fidelity of the specialist models against their respective high-quality references. Pipeline A’s specialist models achieved perfect scores (1.000) as their outputs were used to define the modality-specific references. For image generation, Pipeline C generally outperformed Pipeline B across Image SSIM and Image AverageHash, indicating stronger structural and perceptual similarity to the baseline. However, Pipeline B showed a slightly higher Image HistogramMatch score, suggesting its color distribution was closer to the baseline. This highlights the value of multi-metric evaluation for a nuanced understanding.
@@
Improved Reliability and Safer Deployment The two-part evaluation, enabled by GAICo, is critical for enhancing system reliability. By separating “Plan Coherence” from “Modality Generation Quality”, we can precisely diagnose failures. For instance, the results show that Pipeline B’s poor performance stems from both a weak orchestrator (low scores in Table 1 left) and less-than-optimal specialist models (Table 1 right). This granular insight allows for targeted optimization, improving the orchestrator’s planning logic or replacing a specific specialist model, which is essential for building trustworthy systems and preventing incidents caused by underperforming components. This directly contributes to the goal of producing better AI solutions by enabling a more rigorous and efficient development process.
@@
Table 2: Summary of GAICo Example Notebooks. Those referenced in this paper are emphasized. All available at:
@@ Conclusion
GAICo addresses a critical gap in GenAI evaluation by providing the first unified, extensible, and deployed framework for reproducible assessment across diverse data modalities. Recent studies have found that even with access to the original source code, many ML experiments cannot be reproduced, mainly due to missing details in documentation and inconsistent evaluation procedures (lopresti2021reproducibility). GAICo replaces this current landscape of ad-hoc, non-comparable scripts with a standardized approach, integrating metrics for text, structured data (planning, time-series), and multimedia. This standardization is GAICo’s core novelty; it ensures that once a metric is chosen, its application and reporting are consistent, directly improving the quality and productivity of the evaluation process. As demonstrated in our case study, this unified approach enables nuanced debugging, such as distinguishing orchestrator faults from specialist model failures, which is crucial for enhancing the reliability of complex AI systems.
@@
We now discuss how one may extend the limitations in the future. A key priority is to incorporate metrics for fairness, bias, toxicity, and operational statistics, providing developers with a more holistic toolkit for building safer and more responsible AI systems. GAICo’s extensible design makes this easy. Future work also includes exploring developing interactive dashboards or deeper integration with MLOps platforms (like MLflow (Zaharia_Accelerating_the_Machine_2018)) to enable dynamic exploration and tracking of evaluation results. Lastly, further work would also focus on adding general-purpose comparison capabilities for complex structured data (e.g., JSON diffing, graph similarity) and extending metrics to handle multi-turn conversational outputs.